# IML-Spikeformer: Input-aware Multi-Level Spiking Transformer for Speech Processing

### v1

@@ Abstract
Spiking Neural Networks (SNNs), inspired by biological neural mechanisms, represent a promising neuromorphic computing paradigm that offers energy-efficient alternatives to traditional Artificial Neural Networks (ANNs). Despite proven effectiveness, SNN architectures have struggled to achieve competitive performance on large-scale speech processing task. Two key challenges hinder progress: (1) the high computational overhead during training caused by multi-timestep spike firing, and (2) the absence of large-scale SNN architectures tailored to speech processing tasks. To overcome the issues, we introduce Input-aware Multi-Level Spikeformer, i.e. IML-Spikeformer, a spiking Transformer architecture specifically designed for large-scale speech processing. Central to our design is the Input-aware Multi-Level Spike (IMLS) mechanism, which simulate multi-timestep spike firing within a single timestep using an adaptive, input-aware thresholding scheme. IML-Spikeformer further integrates a Reparameterized Spiking Self-Attention (RepSSA) module with a Hierarchical Decay Mask (HDM), forming the HD-RepSSA module. This module enhances the precision of attention maps and enables modeling of multi-scale temporal dependencies in speech signals. Experiments demonstrate that IML-Spikeformer achieves word error rates of 6.0% on AiShell-1 and 3.4% on Librispeech-960, comparable to conventional ANN transformers while reducing theoretical inference energy consumption by 4.64$\times$ and 4.32$\times$ respectively. IML-Spikeformer marks an advance of scalable SNN architectures for large-scale speech processing in both task performance and energy efficiency.
Index Terms:
Spiking Neural Networks, Neuromorphic Auditory Processing, Speech Recognition, Spiking Transformer
@@ I Introduction
Recent advances in speech processing benefit from large-scale deep learning models, particularly Transformer-based architectures [1]. The high computational cost of such models has motivated the search for energy-efficient alternatives. Spiking Neural Networks (SNNs), which mimic the event-driven information processing of biological neurons, have emerged as a promising solution. Unlike traditional artificial neural networks (ANNs), SNNs emulate the dynamics of biological neurons and utilize spike trains for the representation of information [2]. The event-driven nature of Spiking Neural Networks (SNNs) enables asynchronous computation on neuromorphic hardware [3, 4, 5], where a neuron responds only upon arrival of spike. This sparse activation mechanism significantly reduces power consumption, making SNNs particularly suitable for energy-constrained edge computing applications [6, 7, 8, 9].
@@
Beyond their energy efficiency advantages, SNNs have demonstrated remarkable performance across diverse application domains. Substantive empirical evidence validates their effectiveness in computer vision [10, 11, 12, 13, 14, 15, 16], and natural language processing [17, 18, 19]. In speech processing, SNNs have achieved promising performance in keyword spotting (KWS) [20, 21, 22, 23], acoustic event and sound classification [24, 25], and automatic speech recognition (ASR) [26, 27, 28, 29].
@@
Despite promising results, SNNs are not without issues. For example, the challenges in training strategy[30] and architectures [10] prevent them from scaling up to large networks for complex real-world tasks, such as large-vocabulary automatic speech recognition (ASR) [26]. There have been many attempts to address the issues. By integrating self-attention mechanisms into SNN architecture [13, 14, 16], spiking Transformer models are proposed to enhances model’s sequential modeling capability for long-range and complex temporal dependencies. The spiking Transformer employs multi-timestep firing to increase SNN’s representational power by introducing a performance-efficiency tradeoff. While increasing time window reduces quantization errors [31] and yields substantial performance gains, it also introduces computational complexity that scales linearly with window length, increasing both time and memory requirements [32]. This growing computational overhead ultimately limits the scope of large-scale SNN applications. Moreover, current research on spiking Transformers has predominantly focused on computer vision tasks [13, 14, 16], with their potential for speech processing remaining largely unexplored.
@@
In this work, we aim to develop an spiking Transformer architecture for speech processing tasks with efficient training schemes. To mitigate the computational overhead associated with multi-timestep firing in training, we propose an Input-aware Multi-Level Spike (IMLS) firing mechanism, which facilitates the effective simulation of multi-timestep firing within a single timestep during SNN training, while maintaining the spike-driven computational paradigm in inference. Additionally, IMLS incorporates an input-aware adaptation mechanism that dynamically adjusts neuronal firing thresholds based on the statistical distribution of pre-synaptic inputs, thereby enhancing the representational power of spiking neurons and improving training stability.
@@
The IMLS firing mechanism is the core of the proposed IML-Spikeformer architecture, which features a Hierarchical Decay Re-parameterized Spiking Self-Attention (HD-RepSSA) module. By employing a re-parameterization technique, this module significantly enhances the representational capacity of the spiking attention map while preserving the efficient spike-driven characteristics. Furthermore, drawing inspiration from the neural processing in the auditory system, which exhibits multi-scale temporal dynamics, we introduce a Hierarchical Decay Mask (HDM) to modulate attention maps across various network stages, effectively capturing the multi-scale temporal dependencies that are essential for speech processing tasks. In summary, our contributions in this work are threefold:
@@
  • We introduce the IMLS firing mechanism, which leverages multi-level spike firing to overcome the high training overhead of the multi-timestep firing. The incorporation of an input-aware adaptive neuronal firing threshold that aligns with the pre-synaptic input distribution further enhances the model expressivity and training stability.
@@
  • We design the IML-Spikeformer, a novel spiking Transformer architecture tailored for speech processing. It features the HD-RepSSA module with improved representational capacity with re-parameterized spiking self-attention mechanism. Additionally, by integrating the HDM for hierarchical attention modulation, IML-Spikeformer effectively captures the multi-scale temporal dependencies inherent in speech signals.
@@
  • We demonstrate that the IML-Spikeformer architecture can scale up for in large-scale speech processing tasks, delivering performance comparable to that of ANN-based Transformers while significantly reducing energy consumption in various model sizes. This positions IML-Spikeformer as a promising SNN architecture for energy-efficient speech processing.
@@ II-A Spiking Neural Networks for Speech Processing
SNNs offer the promises for energy efficiency, rapid response, and inherent robustness to noise perturbations. Motivated by these, various SNN architectures have been proposed, such as Spiking Multilayer Perceptrons (SMLP)[27, 33, 34, 35], Spiking Recurrent Neural Networks (SRNN)[36, 37, 6], and Spiking Convolutional Neural Networks (SCNN)[20, 22, 38, 39]. These SNN architectures have demonstrated performance comparable to that of conventional ANNs across various speech processing tasks, including KWS[40, 41, 33, 6, 34], acoustic event and sound classification [42, 43, 44, 24, 45, 46, 47, 48], ASR[26, 27, 28], sound source localization[49, 50, 51], speaker identification [23], and speech enhancement [52, 53, 54, 55].
Despite these advancements, a significant performance gap persists between SNN models and state-of-the-art ANN models in large-scale speech processing tasks. This performance gap primarily stems from previous SNN architectures’ difficulty scaling to larger networks [10], constraining their ability to process complex temporal dependencies and large data scale. Recently, spiking Transformer architecture was a response to the research problem. It shown promising performance on challenging computer vision tasks [13, 56, 29, 14, 16]. However, our preliminary studies indicate that existing spiking Transformer architectures experience significant performance degradation when applied directly to speech processing tasks where there is a need to process speech signals of variable lengths. This is primarily due to the inability of batch normalization in this condition. We consider IML-Spikeformer as an alternative in this paper.
@@ II-B Training Strategy for Large-Scale SNNs
In general, there are two primary approaches for training large-scale SNNs: Spatio-Temporal Backpropagation (STBP)[14, 13, 16, 57] and ANN-to-SNN conversion (ANN2SNN)[58, 29, 59, 60].
STBP explicitly unrolls the network dynamics of SNNs across $T$ timesteps and $L$ layers, applying backpropagation through time (BPTT)[30]. This multi-timestep simulation incurs $O(L\times T)$ memory complexity to store the neuronal states necessary for gradient computation, as well as $O(T)$ time complexity due to its inherently iterative processing. On the other hand, ANN2SNN methods convert a pre-trained ANN into SNN via rate-coding approximation, effectively circumventing the training overhead since the ANN2SNN conversion only need finetuning for few epochs. However, such technique necessitates many timesteps (like 16)[58, 59, 60] to accurately approximate the activations of the ANN with spike firing rate, resulting in high computational costs during inference. Both methods introduce substantial computational overhead—STBP during training and ANN2SNN during inference—making the implementation of large-scale SNNs more expensive.
@@ III-A Spiking Neuron Models
Spiking neuron models are computational abstraction motivated by the understanding of biological neurons. Among these, the Leaky Integrate-and-Fire (LIF) model [2] is widely adopted for constructing large-scale SNNs due to its mathematical tractability. The neuronal dynamics of the LIF model can be described by the following discrete-time formulation:
$v^{l}[t]=\beta\hskip 1.0ptv^{l}[t-1]+x^{l-1}[t],$ (1)
@@
$v^{l}[t]=v^{l}[t]-\theta s^{l}[t],$ (3)
where $t$ denotes the timestep, $\theta$ is the firing threshold. $x^{l-1}[t]$ is the pre-synaptic input from the layer $l-1$, $v^{l}[t]$ and $s^{l}[t]$ denote the membrane potential and the output spike in layer $l$, respectively. Notably, Eqs. 1-3 describe the leakage & integration, firing, and reset processes of the spiking neuron, respectively.
Leakage & Integration: This process defines two essential dynamics within a spiking neuron, including the decay of information according to a leaky factor $0\leq\beta\leq 1$ and the integration of the presynaptic input $x^{l-1}[t]$. When $\beta=1$, the neuron functions as an Integrate-and-Fire (IF) neuron, indicating no information leakage between timesteps; otherwise, it operates as a LIF neuron.
Firing: The function $\mathcal{F}(\cdot)$ represents the spike firing mechanism, and $s^{l}[t]$ denotes the output spike. When the membrane potential $v^{l}[t]$ surpasses the firing threshold $\theta$, an output spike is generated followed by the membrane potential reset.
Reset: The membrane potential resets immediately after neuron firing by subtracting $\theta$ from the membrane potential $v^{l}[t]$, which is known as a soft reset.
@@ IV-A Multi-Level Spike Firing
$s^{M}$ =$=\begin{cases}0,&\quad v[1]<\theta\\[6.0pt] n,&\quad(n-1)\theta\leq v[1]<n\theta\\[6.0pt] T,&\quad(T-1)\theta\leq v[1]\end{cases}$ (7)
@@
[…]
$=\frac{\partial\mathrm{clip}(\frac{v[1]}{\theta},0,T)}{\partial v[1]}=\begin{cases}\frac{1}{\theta},&\text{if }0\leq v[1]\leq\theta T\\ 0.&\text{Otherwise}\end{cases}$
@@ IV-B Input-aware Multi-Level Spike Firing
For stable spike firing, we propose Input-aware Multi-Level Spike (IMLS) firing mechanism, which extend the MLS firing mechanism with adaptive thresholds that automatically adjust to pre-synaptic input, allowing our model to maintain stable spike firing patterns even without explicit normalization like BN. The IMLS firing mechanism incorporates a channel-wise adaptive scaling factor $\lambda\in\mathbb{R}^{C}$ for the firing threshold $\theta$. Unlike fixed thresholds, the adaptive threshold $\lambda\theta$ dynamically adjusts in response to the statistical distribution of pre-synaptic inputs $x[1]$.
@@ V-A Re-parametrized Spiking Self-Attention
$\mathbf{A}$ =$=\begin{cases}XW_{Q}W_{K}^{T}X^{T},&\text{Training}\\ XW_{QK}X^{T},&\text{Inference}\end{cases}$ (12)
@@ V-B Hierarchical Decay Mask
where $\phi(l)$ represents the layer-specific decay function, formulated as $\phi(l)=1-2^{-5-l}$. The term $|i-j|$ denotes the absolute positional distance between the tokens. The decay factor increases as the layer depth increases, enforcing a structured attention modulation strategy: shallow layers apply more pronounced attenuation to prioritize local interactions, while deeper layers retain decay values near 1, preserving long-range dependencies crucial for capturing global context.
The HDM can be integrated into the proposed RepSSA through element-wise multiplication with the attention map, creating a biologically-inspired attention mechanism that inherently captures multi-scale temporal dependencies. This integration forms the core component of IML-Spikeformer, which termed HD-RepSSA, formulated as:
[…]
@@ VI-A Task Configurations
To assess the effectiveness and efficiency of the proposed IML-Spikeformer model, we conducted extensive experiments on large-scale speech processing tasks where SNNs used to achieve inferior performance to ANNs. In all the experiments, IML-Spikeformer is implemented as a direct substitute for ANN transformer.
ASR: The proposed IML-Spikeformer was evaluated on two public ASR databases, AiShell-1 [70] that consists of 170 hours of Mandarin speech data, and LibriSpeech-960 [71] of approximately 960 hours of English audiobook data. The experiment pipelines follow the established ESPnet ASR recipes [72] for data preparation, model training, and evaluation protocols. We use the ANN Transformers, IML-Spikeformer and other baseline models (detailed in Section VI-B) as the encoder for ASR experiments, with a fixed 6-layer Transformer as the decoder without a language model in all experiments to ensure a fair comparison. In data pre-processing, an 80-dimensional log Mel filterbank is used to represent the acoustic features. Data augmentation techniques were systematically applied for improved robustness. Specifically SpecAugment [73] was adopted for frequency and time masking, while speed perturbation was adopted for increased variability. The input features were subsequently downsampled through two consecutive 3$\times$3 convolutional layers with a stride of 2, reducing the temporal dimensionality of original speech inputs.

### v2

@@ Abstract
[…]
@@ I Introduction
Recent advances in speech processing benefit from large-scale deep learning models, particularly Transformer-based architectures [1]. The high computational cost of such models has motivated the search for energy-efficient alternatives. Spiking Neural Networks (SNNs), which mimic the event-driven information processing of biological neurons, have emerged as a promising solution. Unlike traditional artificial neural networks (ANNs), SNNs emulate the dynamics of biological neurons and utilize spike trains for the representation of information [2]. The event-driven nature of Spiking Neural Networks (SNNs) enables asynchronous computation on neuromorphic hardware [3, 4, 5], where a neuron responds only upon arrival of a spike. This sparse activation mechanism significantly reduces power consumption, making SNNs particularly suitable for energy-constrained edge computing applications [6, 7].
@@
Beyond their energy efficiency advantages, SNNs have demonstrated remarkable performance across diverse application domains. Substantive empirical evidence validates their effectiveness in computer vision [8, 9, 10, 11], and natural language processing [12, 13]. In speech processing, SNNs have achieved promising performance in keyword spotting (KWS) [14, 15, 16], acoustic event and sound classification [17, 18], and automatic speech recognition (ASR) [19, 20, 21, 22].
@@
Despite promising results, SNNs are not without issues. For example, the challenges in training strategy[23, 24] and architectures [8, 25] prevent them from scaling up to large networks for complex real-world tasks, such as large-vocabulary ASR [19]. There have been many attempts to address the issues. By integrating self-attention mechanisms into SNN architecture [9, 26, 11], spiking Transformer models are proposed to enhances model’s sequential modeling capability for long-range and complex temporal dependencies. The spiking Transformer employs multi-timestep firing to increase SNN’s representational power by introducing a performance-efficiency tradeoff.
While several methods [27, 28, 29, 30] are proposed to reduce the information loss in spike firing [31] and yields substantial performance gains with increase time windows, it also introduces computational complexity that scales linearly with window length, increasing both time and memory requirements [32]. This growing computational overhead ultimately limits the scope of large-scale SNN applications [33, 34]. Moreover, current research on spiking Transformers has predominantly focused on computer vision tasks [9, 26, 11], with their potential for speech processing remaining largely unexplored.
@@
In this work, we aim to develop a spiking Transformer architecture for speech processing tasks with efficient training schemes. To mitigate the computational overhead of multi-timestep firing during training, we propose an Input-aware Multi-Level Spike (IMLS) firing mechanism. IMLS enables the simulation of multi-timestep dynamics within a single timestep, thereby reducing training costs while retaining the spike-driven computational paradigm at inference. Moreover, IMLS introduces an input-aware adaptation strategy that adjusts neuronal firing thresholds according to the statistical distribution of pre-synaptic inputs, which improves both representational expressivity and training stability.
@@
Building on IMLS, we design the IML-Spikeformer, a spiking Transformer tailored for speech processing. Its central component is the Hierarchical Decay Re-parameterized Spiking Self-Attention (HD-RepSSA) module, which enhances the representational capacity of the spiking attention map via a re-parameterization scheme while preserving spike-driven efficiency. Inspired by the auditory system’s multi-scale temporal dynamics, we further propose a Hierarchical Decay Mask (HDM) that modulates attention maps across network layers, enabling the model to capture multi-scale temporal dependencies crucial for speech tasks.
Our main contributions are summarized as follows:
@@
  • IMLS mechanism: We introduce a multi-level spike firing strategy that reduces the training cost of multi-timestep firing, while an input-aware adaptive threshold aligned with pre-synaptic input statistics enhances model expressivity and stabilizes training.
@@
  • IML-Spikeformer architecture: We propose a novel spiking Transformer featuring the HD-RepSSA module for re-parameterized spiking self-attention and the HDM for hierarchical temporal modulation, enabling effective modeling of speech signals with multi-scale dynamics.
@@
  • Scalability and efficiency: We demonstrate that IML-Spikeformer scales to large speech processing tasks, achieving performance comparable to ANN-based Transformers while substantially reducing energy consumption across model sizes, highlighting its promise for energy-efficient speech processing.
@@ II-A Spiking Neural Networks for Speech Processing
SNNs offer the promises of energy efficiency, rapid response, and inherent robustness to noise perturbations. Motivated by these, various SNN architectures have been proposed, such as Spiking Multilayer Perceptrons (SMLP)[20, 35], Spiking Recurrent Neural Networks (SRNN)[36, 37], and Spiking Convolutional Neural Networks (SCNN)[38, 15]. These SNN architectures have demonstrated performance comparable to that of conventional ANNs across various speech processing tasks, including KWS[39, 40], acoustic event and sound classification [17, 41], ASR[19, 20, 21], sound source localization[42, 43], speaker identification [16], and speech enhancement [44, 45, 46].
Despite these advancements, a significant performance gap persists between SNN models and state-of-the-art ANN models in large-scale speech processing tasks. This performance gap primarily stems from previous SNN architectures’ difficulty scaling to larger networks [8], constraining their ability to process complex temporal dependencies and large data scale. Recently, spiking Transformer architecture was a response to the research problem. It shown promising performance on challenging computer vision tasks [9, 22, 11]. However, our preliminary studies indicate that existing spiking Transformer architectures experience significant performance degradation when applied directly to speech processing tasks where there is a need to process speech signals of variable lengths. This is primarily due to the inability of batch normalization in this condition. We consider IML-Spikeformer as an alternative in this paper.
@@ II-B Training Strategy for Large-Scale SNNs
In general, there are two primary approaches for training large-scale SNNs: Spatio-Temporal Backpropagation (STBP)[26, 9, 47] and ANN-to-SNN conversion (ANN2SNN)[48, 22, 49].
STBP explicitly unrolls the network dynamics of SNNs across $T$ timesteps and $L$ layers, applying backpropagation through time (BPTT)[23]. This multi-timestep simulation incurs $O(L\times T)$ memory complexity to store the neuronal states necessary for gradient computation, as well as $O(T)$ time complexity due to its inherently iterative processing. On the other hand, ANN2SNN methods convert a pre-trained ANN into SNN via rate-coding approximation, effectively circumventing the training overhead since the ANN2SNN conversion only need finetuning for few epochs. However, such technique necessitates many timesteps (like 16)[48, 49] to accurately approximate the activations of the ANN with spike firing rate, resulting in high computational costs during inference. Both methods introduce substantial computational overhead—STBP during training and ANN2SNN during inference—making the implementation of large-scale SNNs more expensive.
@@ III-A Spiking Neuron Models
Spiking neuron models are computational abstractions motivated by the understanding of biological neurons, acting as the basic computational units of SNNs. Among these, the Leaky Integrate-and-Fire (LIF) model [2] is widely adopted for constructing large-scale SNNs due to its mathematical tractability. The neuronal dynamics of the LIF model can be described by the following discrete-time formulation:
$v^{l}[t]=\beta\hskip 1.0ptv^{l}[t-1]+x^{l-1}[t]-\theta s^{l}[t-1],$ (1)
@@
where $t$ denotes the timestep, $\theta$ is the firing threshold, $x^{l-1}[t]$ is the pre-synaptic input from layer l1l-1, and $v^{l}[t]$ and $s^{l}[t]$ denote the membrane potential and output spike in layer $l$, respectively. Equations 1-2 describe three fundamental processes of the spiking neuron: leakage & integration, reset, and firing.
Leakage & Integration: This process defines two essential dynamics within a spiking neuron: the decay of information according to a leaky factor $0\leq\beta\leq 1$ and the integration of the pre-synaptic input $x^{l-1}[t]$. When $\beta=1$, the neuron functions as an Integrate-and-Fire (IF) neuron with no information leakage between timesteps; otherwise, it operates as a LIF neuron.
Reset: After integration, the membrane potential undergoes soft reset by subtracting $\theta$ from neurons that fired in the previous timestep, as represented by the last term in Eq. 1.
Firing: The function $\mathcal{F}(\cdot)$ represents the spike firing mechanism. When the membrane potential $v^{l}[t]$ surpasses the firing threshold $\theta$, an output spike $s^{l}[t]=1$ is generated; otherwise, $s^{l}[t]=0$.
@@ IV-A Multi-Level Spike Firing
[…]
@@
where $n\in{1,\dots,T-1}$, $\lfloor\cdot\rfloor$ denotes the floor function, and $v[1]$ represents the membrane potential resulting from the integration of pre-synaptic input $x[1]$ at the initial timestep.
This multi-level spike representation in MLS shares similarities with burst coding [50]. Recent implementations [51, 52, 53] also employ spike magnitude to encode enhanced neural activity beyond binary spikes. However, the fundamental difference lies in their temporal dynamics: while burst coding only considers the number of spikes within a time interval, MLS establishes a mathematical equivalence between multi-timestep IF neuron firing and multi-level spike values. This equivalence (shown in Fig. 1(a)) enables seamless conversion from integer spike computing to binary spike-driven computing during inference.
Since the $\lfloor\cdot\rfloor$ is non-differentiable at the boundaries, we leverage the Generalized Straight-Through Estimator (G-STE) [54] to approximate the gradient, yielding:
@@ IV-B Input-aware Multi-Level Spike Firing
For stable spike firing, we propose Input-aware Multi-Level Spike (IMLS) firing mechanism, which extends the MLS firing mechanism with adaptive thresholds that automatically adjust to pre-synaptic input, allowing our model to maintain stable spike firing patterns even without explicit normalization like BN. The IMLS firing mechanism incorporates a channel-wise adaptive scaling factor $\lambda\in\mathbb{R}^{C}$ for the firing threshold $\theta$. Unlike fixed thresholds, the adaptive threshold $\lambda\theta$ dynamically adjusts in response to the statistical distribution of pre-synaptic inputs $x[1]$.
@@ V-A Re-parametrized Spiking Self-Attention
[…]
@@ V-B Hierarchical Decay Mask
where $\phi(l)$ represents the layer-specific decay function, formulated as $\phi(l)=1-2^{-5-l}$. The term |ij||i-j| denotes the absolute positional distance between the tokens. The decay factor increases as the layer depth increases, enforcing a structured attention modulation strategy: shallow layers apply more pronounced attenuation to prioritize local interactions, while deeper layers retain decay values near 1, preserving long-range dependencies crucial for capturing global context.
The HDM can be integrated into the proposed RepSSA through element-wise multiplication with the attention map, creating a biologically-inspired attention mechanism that inherently captures multi-scale temporal dependencies. This integration forms the core component of IML-Spikeformer, which is termed HD-RepSSA, formulated as:
$\text{HD-RepSSA}(\mathbf{A},\mathbf{V_{s}})=\text{RepSSA}(\mathbf{A^{\prime}},\mathbf{V_{s}}),\hskip 3.0pt\mathbf{A^{\prime}}=\mathbf{A}\odot\mathbf{H}.$ (14)
@@ VI-A Task Configurations
To assess the effectiveness and efficiency of the proposed IML-Spikeformer model, we conducted extensive experiments on large-scale speech processing tasks where SNNs historically achieved inferior performance compared to ANNs. In all experiments, IML-Spikeformer is implemented as a direct substitute for ANN Transformer.
ASR: The proposed IML-Spikeformer was evaluated on two public ASR databases: AiShell-1 [62], consisting of 170 hours of Mandarin speech data, and LibriSpeech-960 [63], comprising approximately 960 hours of English audiobook data. The experimental pipelines follow the established ESPnet ASR recipes [64] for data preparation, model training, and evaluation protocols. We employ ANN Transformers, IML-Spikeformer, and other baseline models (detailed in Section VI-B) as encoders, with a fixed 6-layer Transformer decoder without language model across all experiments to ensure fair comparison.
For evaluation, we assess model performance using Character Error Rate (CER) on AiShell-1’s development and test sets (labeled as ”dev” and ”test” in Table II). For LibriSpeech-960, we evaluate using Word Error Rate (WER) on both clean and other subsets of the development and test sets. The clean subset contains high-quality recordings with minimal background noise and clear pronunciation, while the other subset includes more challenging audio with background noise, varied recording conditions, and diverse speaker accents, providing comprehensive assessment across different acoustic conditions.
In data preprocessing, 80-dimensional log Mel filterbank features represent the acoustic inputs. Data augmentation techniques are systematically applied for improved robustness: SpecAugment [65] for frequency and time masking, and speed perturbation for temporal variability. The input features are subsequently downsampled through two consecutive 3×3 convolutional layers with stride 2, reducing the temporal dimensionality of the original speech inputs.